# Deformation dynamics of an oil droplet into a crescent shape during the intermittent motion

### v2

@@ I Introduction
For the self-propelled motion of a solid object, its shape determines the motion. On the other hand, the feedback between motion and shape can be anticipated for self-propelled soft objects like droplets because motion can modify the object shape, and the shape affects its motion. The interactions between motion and shape of self-propelled droplets were studied in a number of papers[23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35]. The mechanism of the coupling between motion and deformation is an essential problem in self-propulsion in non-equilibrium systems. There have been several theoretical reports in which generic models for the coupling between motion and deformation by considering the symmetric properties were discussed[36, 37, 38]. These generic models are also applicable to the biological cell motion with deformation[39, 40]. However, these studies were concerned with small magnitudes of deformation from a circle. As far as the authors know, there has been no universal theoretical description that can be applied to large deformation.
@@
In the presented study, we investigated the mechanism of deformation of a paraffin droplet containing camphor and oil red O, in which intermittent motion with dynamic deformation can be observed depending on the concentrations of camphor and oil red O and the elapsed time from putting the droplet on the water surface. We mainly focused on the change between the crescent and circular shapes observed during the intermittent motion. We introduced the quantities that allow us to measure the degree of deformation and evaluated them by image processing of the experimental results. Then, we discussed the relationship between these quantities and the droplet speed. We also measured the interfacial tension between water and paraffin containing oil red O and proposed the deformation mechanism towards a crescent shape coupled with intermittent motion.
@@ III.1 Spontaneous motion of a droplet
A $20~{}\mu\mathrm{L}$ droplet of paraffin with dissolved camphor (concentration $0.06~{}\mathrm{M}$) and oil red O (concentration $1.0~{}\mathrm{g/L}$) exhibited three different types of motion depending on the time $t$ elapsed from the moment when the droplet was placed on the water surface. At the beginning of the experiment, the droplet showed self-propelled motion along a straight line perturbed by reflections with the dish wall, as illustrated in Fig. 1(a-1). Approximately at $t=30~{}\mathrm{s}$, the droplet began to rotate along the wall (Fig. 1(a-2)). Then, at $t\sim 2000~{}\mathrm{s}$, the droplet motion switched to intermittent one (Fig. 1(a-3)). At this stage, short intervals of time when a droplet moved (less than $10~{}\mathrm{s}$) were separated by periods when it was standing. The droplet had a circular shape when it did not move, and it became elongated and exhibited deformation into a crescent shape during intermittent motion.
@@
Figure 3: Dynamics of a single phase of the intermittent motion of a $20~{}\mu\mathrm{L}$ paraffin droplet with camphor and oil red O. The concentrations of camphor and oil red O were $0.06~{}\mathrm{M}$ and $1.0~{}\mathrm{g/L}$, respectively. (a) The speed $v$ of the droplet COM, the aspect ratio $R$, and the asymmetric parameter $A$ as functions of $\Delta t$. The time $\Delta t=0$ corresponds to when $v$ has the maximum value. (b) A sequence of frames showing droplet location and deformation during the phase of intermittent motion described in (a).
Figure 3(a) shows the speed $v(t)$ of the droplet center $(\bm{r}_{\mathrm{c}}(t))$ and the deformations $R(t)$ and $A(t)$ as functions of time during a single phase of intermittent motion illustrated in Fig. 3(b). The time $\Delta t=0$ corresponds to the moment when $v(t)$ has the maximum value. As shown in Fig. 3(a), the deformations began almost simultaneously when motion started. The speed $v(t)$ reached its maximum before the maxima of $R(t)$ and $A(t)$, which means that deformations still developed after the droplet started to slow down. This type of behavior was reproduced in many observations of intermittent motion, as shown in Fig. 4, in which the scatter plot of the time $\Delta t_{R}$ for the maximum of $R(t)$ and $\Delta t_{A}$ for the maximum of $A(t)$ is displayed. In almost all analyzed cases, both $\Delta t_{R}$ and $\Delta t_{A}$ were positive, and $\Delta t_{A}$ was slightly greater than $\Delta t_{R}$.
@@
Next, we investigated the dependence of the droplet deformation on the oil red O concentration. We performed experiments at oil red O concentrations in the range of 0.50.50.50.51.$1.1~{}\mathrm{g/L}$. The maximum values of $R(t)$ and $A(t)$ ($R_{\mathrm{max}}$ and $A_{\mathrm{max}}$), increased for oil red O concentrations larger than $0.8~{}\mathrm{g/L}$ as shown in Fig. 5.
@@ III.2 Interfacial tension between oil and water
where $\gamma_{\mathrm{ow}}^{(0)}$ is the interfacial tension at the initial stage, $\Delta\gamma_{\mathrm{ow}}$ is the difference in interfacial tension between the initial and final stages, and $\tau$ is the characteristic time for the decrease in interfacial tension. We plotted $\gamma_{\mathrm{ow}}^{(0)}$, $\Delta\gamma_{\mathrm{ow}}$ and $\tau$ against oil red O concentration in Fig. 7. $\gamma_{\mathrm{ow}}^{(0)}$ and $\Delta\gamma_{\mathrm{ow}}$ were decreasing and increasing functions of oil red O concentration, respectively. In the case when the decrease in interfacial tension was clearly observed (oil red O concentrations: $0.5$, $0.8$, and $1.0~{}\mathrm{g/L}$), $\tau$ was approximately equal to $800~{}\mathrm{s}$. In addition, we performed the same measurement for the oil droplet that contained both oil red O ($1.0~{}\mathrm{g/L}$) and camphor ($0.06~{}\mathrm{M}$). The decrease in interfacial tension was also observed, and $\gamma_{\mathrm{ow}}^{(0)}$ was lower compared to paraffin droplet with oil red O ($1.0~{}\mathrm{g/L}$) but without camphor.
Figure 6: Time evolution of the interfacial tension $\gamma_{\mathrm{ow}}(t)$ between a paraffin droplet and pure water measured by the pendant drop method. The time $t=0$ corresponds to when a droplet was formed. Marker colors code the oil red O concentrations: $0~{}\mathrm{g/L}$ black, $0.1~{}\mathrm{g/L}$ yellow, $0.5~{}\mathrm{g/L}$ orange, $0.8~{}\mathrm{g/L}$ red, $1.0~{}\mathrm{g/L}$ dark red. For each concentration, the average values of four experiments are plotted. The error bars indicate the standard deviation of the results. $\gamma_{\mathrm{ow}}(t)$ between pure water and paraffin droplet that contained both oil red O ($1.0~{}\mathrm{g/L}$) and camphor ($0.06~{}\mathrm{M}$) is marked by
blue squares.
Figure 7: Parameters fitting $\gamma_{\mathrm{ow}}(t)$ according to Eq. (1) for different concentrations of oil red O (cf. Fig. 6). (a) Interfacial tension on the initial stage $\gamma_{\mathrm{ow}}^{(0)}$. (b) Difference in interfacial tension $\Delta\gamma_{\mathrm{ow}}$ between the initial and final stages. (c) Characteristic time $\tau$ for the decrease in interfacial tension. The data for a paraffin droplet that contained both oil red O ($1.0~{}\mathrm{g/L}$) and camphor ($0.06~{}\mathrm{M}$) are shown with blue squares. The average values for four measurements are shown. The error bars indicate their standard deviations.
@@ IV.1 Analysis of free energy related to droplet deformation
To better understand the droplet deformation phenomenon, we evaluated two related dimensionless numbers. First, the capillary number $\mathrm{Ca}$ is the ratio of the viscous force and interfacial tension, $\mathrm{Ca}=\mu v/\gamma_{\mathrm{ow}}$, where $\mu$ is the viscosity of the liquid, $v$ is the characteristic velocity, and $\gamma_{\mathrm{ow}}$ is the characteristic value for oil/water interfacial tension. By substituting $\mu\simeq 10^{-3}~{}\mathrm{Pa\,s}$, $v\simeq 10^{-2}~{}\mathrm{m/s}$, and $\gamma_{\mathrm{ow}}\simeq 60\times 10^{-3}~{}\mathrm{N/m}$, we obtain $\mathrm{Ca}\simeq 2\times 10^{-4}\ll 1$. This means that during an intermittent motion, the relative effect of viscous drag forces is much smaller than that of surface tension.
Next, the Weber number $\mathrm{We}$ is the relative importance of the fluid inertia compared to its surface tension. $\mathrm{We}=\rho rv^{2}/\gamma_{\mathrm{ow}}$, where $\rho$ is the mass density of the liquid and $r$ is the characteristic length. For $\rho\simeq 10^{3}~{}\mathrm{kg/m^{3}}$, $r\simeq 10^{-2}~{}\mathrm{m}$ and other parameter values listed above, we obtained $\mathrm{We}\simeq 2\times 10^{-4}\ll 1$. This means that the inertial effects are negligible during an intermittent motion. Considering both the dimensionless numbers, we concluded that the surface tension difference around the droplet plays the most important role in droplet deformation.
@@
Here, $S=\gamma_{\mathrm{aw}}-(\gamma_{\mathrm{ow}}+\gamma_{\mathrm{ao}})$ is the spreading coefficient, $\gamma_{\mathrm{aw}}$ and $\gamma_{\mathrm{ao}}$ are the interfacial tensions between air and water and between air and oil, respectively, $\Sigma$ is the area of the droplet, and $\Delta\rho_{\mathrm{ao}}$ and $\Delta\rho_{\mathrm{ow}}$ are the density differences between air and oil and that between oil and water, respectively, as shown in Fig. 8. In Eq. (2), $e_{\mathrm{ao}}$ and $e_{\mathrm{ow}}$ are the differences in height between the water level at infinity and the air/oil interface, and that between the water level at infinity and the oil/water interface, respectively. For the following analyses, we neglected the second term in Eq. (2) as it is much smaller than the first term since the thickness of a droplet is small.
Figure 8: Schematic illustration of an oil droplet floating on the water surface. The triple line matches $x=z=0$. The functions $z=\zeta_{\mathrm{ao}}(x)$ and $z=-\zeta_{\mathrm{ow}}(x)$ indicate the interface between air and oil and that between oil and water, respectively. $\theta_{\mathrm{ao}}(x)$ and $\theta_{\mathrm{ow}}(x)$ are the slopes of them, i.e., $\tan\theta_{\mathrm{ao}}(x)=\mathrm{d}\zeta_{\mathrm{ao}}/\mathrm{d}x$ and $\tan\theta_{\mathrm{ow}}(x)=\mathrm{d}\zeta_{\mathrm{ow}}/\mathrm{d}x$. $e_{\mathrm{ao}}$ and $e_{\mathrm{ow}}$ are the differences in height between the water level at infinity and the air/oil interface, and that between the water level at infinity and the oil/water interface, respectively.
@@
$\mathcal{T}=$ $\frac{4}{3}\gamma_{\mathrm{ao}}\kappa_{\mathrm{ao}}^{-1}\left(1-\cos^{2}\frac{\theta_{\mathrm{ao}}(x=0)}{2}\right)$ (3)
$+\frac{4}{3}\gamma_{\mathrm{ow}}\kappa_{\mathrm{ow}}^{-1}\left(1-\cos^{2}\frac{\theta_{\mathrm{ow}}(x=0)}{2}\right),$
where $\kappa_{\mathrm{ao}}^{-1}$ and $\kappa_{\mathrm{ow}}^{-1}$ are capillary lengths of air/oil and oil/water interfaces, respectively. $\theta_{\mathrm{ao}}(x=0)$ is the angle between the air/oil interface and the $x$-axis, and $\theta_{\mathrm{ow}}(x=0)$ is that between the oil/water interface and the $x$-axis. These values can be calculated from the balance at the triple line. That is, by assuming the water surface outside of the droplet is completely flat, we obtain
$\cos\theta_{\mathrm{ao}}(x=0)$ $=\frac{\gamma_{\mathrm{aw}}^{2}-\gamma_{\mathrm{ow}}^{2}+\gamma_{\mathrm{ao}}^{2}}{2\gamma_{\mathrm{aw}}\gamma_{\mathrm{ao}}},$ (4)
$\cos\theta_{\mathrm{ow}}(x=0)$ $=\frac{\gamma_{\mathrm{aw}}^{2}-\gamma_{\mathrm{ao}}^{2}+\gamma_{\mathrm{ow}}^{2}}{2\gamma_{\mathrm{aw}}\gamma_{\mathrm{ow}}}.$ (5)
The detail in the derivation is presented in the Appendix A.3.
@@
To support self-motion with deformation, the energy gain due to the surface tension gradient should exceed the free energy increase due to the droplet deformation depending on the interfacial tension. In our system, the quasi-elastic laser scattering (QELS) measurements shown in Appendix B suggested that the value of $\delta\gamma_{\mathrm{aw}}$ does not exceed $0.3~{}\mathrm{mN/m}$, though we were not able to measure this value directly. For the following calculations, we used $\delta\gamma_{\mathrm{aw}}=0.25~{}\mathrm{mN/m}$. During the intermittent motion shown in Fig. 3, we observed $\Delta\Sigma=3.3~{}\mathrm{mm}^{2}$, $\Delta l=17.2~{}\mathrm{mm}$ and $\sigma=1288.5~{}\mathrm{mm}^{2}$. For these parameters and values of interfacial tensions measured by the Wilhelmy method: $\gamma_{\mathrm{ao}}=31.6~{}\mathrm{mN/m}$ and $\gamma_{\mathrm{aw}}=73.5~{}\mathrm{mN/m}$, the values of $\Delta E_{\Sigma}=S\Delta\Sigma$, $\Delta E_{l}=\mathcal{T}\Delta l$ and $W$ are plotted as functions of $\gamma_{\mathrm{ow}}$ in Fig. 9(a). For these realistic values of parameters ($\gamma_{\mathrm{ow}}\sim 50~{}\mathrm{mN/m}$), $\Delta E_{\Sigma}+\Delta E_{l}$, and $W$ are of the same order, which implies that the deformation of a droplet can occur. $\Delta E_{\Sigma}(\gamma_{\mathrm{ow}})$ and $\Delta E_{l}(\gamma_{\mathrm{ow}})$ are increasing functions of interfacial tension $\gamma_{\mathrm{ow}}$, while $W$ does not depend on it. The increase in $\Delta E_{l}(\gamma_{\mathrm{ow}})$ is faster than that in $\Delta E_{\Sigma}(\gamma_{\mathrm{ow}})$, which means that the line energy is the most important for droplet deformation. Figure 9(b) shows the deformation energy normalized by the work exerted on the droplet. For the selected parameter values, the spontaneous droplet deformation can occur when $\Delta E/W<1$, which happens if $\gamma_{\mathrm{ow}}\lesssim 47~{}\mathrm{mN/m}.$ Having in mind that $\gamma_{\mathrm{ow}}$ is a decreasing function of oil red O concentration (cf. Fig. 6), we can see why droplet deformation occurs for droplets with a higher concentration of oil red O. The estimated threshold for deformation is a bit lower than expected from the results discussed in the previous chapter. There could be many reasons for this. First, the estimation is based on a simple model. Second, the local inhomogeneities in oil red O concentration, especially those near the droplet boundary, can trigger local shape changes. And finally, there may be possible errors in the measurements of interfacial tension. For example, the threshold should increase for the lower value of $\gamma_{\mathrm{ao}}$.
@@
Figure 9: Major forms of energy related to droplet deformation as functions of the interfacial tension between oil and water. (a) Plots of $\Delta E_{\Sigma}$ (red), $\Delta E_{l}$ (orange) and $W$ (blue), respectively. (b) Deformation energy normalized by the work exerted on the droplet, $\Delta E/W$ against $\gamma_{\mathrm{ow}}$.

### v3

# Deformation dynamics of an oil droplet into a crescent shape during intermittent motion

@@ I Introduction
For the self-propelled motion of a solid object, its shape determines the motion. In the case of a soft, deformable, self-propelled object such as an oil droplet containing camphor, the motion is also initiated and supported by asymmetric concentration of surface active molecules around. However, for such an object, the feedback between motion and shape can be anticipated because the motion can modify the object shape, and the shape affects its motion. The interactions between motion and shape of self-propelled droplets were studied in a number of papers[23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35]. The mechanism of the coupling between motion and deformation is an essential problem in self-propulsion in non-equilibrium systems. There have been several theoretical reports in which generic models for the coupling between motion and deformation by considering the symmetric properties were discussed[36, 37, 38]. These generic models are also applicable to the biological cell motion with deformation[39, 40]. However, these studies were concerned with small magnitudes of deformation from a circle. As far as the authors know, there has been no universal theoretical description that can be applied to large deformation.
@@
In the presented study, we investigated the mechanism of deformation of a paraffin droplet containing camphor and oil red O, in which intermittent motion with dynamic deformation can be observed depending on the concentrations of camphor and oil red O and the elapsed time from putting the droplet on the water surface. We mainly focused on the change between the crescent and circular shapes observed during the intermittent motion. We introduced the quantities that allow us to measure the degree of deformation and evaluated them by image processing of the experimental results. Then, we discussed the relationship between these quantities and the droplet speed. We also measured the interfacial tension between water and paraffin containing oil red O and explained the deformation mechanism towards a crescent shape coupled with intermittent motion.
@@ III.1 Spontaneous motion of a droplet
The motion of a $20~{}\mu\mathrm{L}$ droplet of paraffin with dissolved camphor (concentration $0.06~{}\mathrm{M}$) and oil red O (concentration $1.0~{}\mathrm{g/L}$) clearly demonstrated the transient character of the phenomenon. Three different types of motion could be easily distinguished for different times $t$ elapsed from the moment when the droplet was placed on the water surface. At the beginning of the experiment, the droplet showed self-propelled motion along a straight line perturbed by reflections with the dish wall, as illustrated in Fig. 1(a-1). Approximately at $t=30~{}\mathrm{s}$, the droplet began to rotate along the wall (Fig. 1(a-2)). Such type of motion was observed for over $1500~{}\mathrm{s}$. The droplet speed began to oscillate with an increasing amplitude and a decreasing average as illustrated in Fig. 1(b). Then, at $t\sim 2000~{}\mathrm{s}$, the droplet motion switched to intermittent one (Fig. 1(a-3)). At this stage, short intervals of time when a droplet moved (less than $10~{}\mathrm{s}$) were separated by periods when it was standing. The droplet had a circular shape when it did not move, and it became elongated and exhibited deformation into a crescent shape during intermittent motion. We observed a monotonic decrease in the peak speed as a function of time. It can be anticipated that the changes in the character of motion are related to the rate of camphor release, which decreases the water surface tension around it. However, as we discuss below, the slow modification of interfacial tension of the oil-water interface resulting from the presence of dye molecules near the surface is another factor leading to intermittent motion.
@@
Figure 3: Dynamics of a single phase of the intermittent motion of a $20~{}\mu\mathrm{L}$ paraffin droplet with camphor and oil red O. The concentrations of camphor and oil red O were $0.06~{}\mathrm{M}$ and $1.0~{}\mathrm{g/L}$, respectively. (a) The speed $v$ of the droplet COM, the aspect ratio $R$, and the asymmetric parameter $A$ as functions of $\Delta t$. The time $\Delta t=0$ corresponds to when $v$ has the maximum value. (b) A sequence of frames showing droplet location and deformation during the phase of intermittent motion analyzed in (a).
Figure 3(a) shows the speed $v(t)$ of the droplet center $(\bm{r}_{\mathrm{c}}(t))$ and the deformations $R(t)$ and $A(t)$ as functions of time during a single phase of intermittent motion illustrated in Fig. 3(b). The time $\Delta t=0$ corresponds to the moment when $v(t)$ has the maximum value. As shown in Fig. 3(a), the deformations began almost simultaneously when motion started. The speed $v(t)$ reached its maximum before the maxima of $R(t)$ and $A(t)$, which means that deformations still developed after the droplet started to slow down. The comparison between $v(t)$, $R(t)$ and $A(t)$ shows that the maximum acceleration appeared when the values of both the parameters describing deformation were around $50\%$ of their maximum.
The behavior illustrated in Fig. 3(a) was reproduced in many observations of intermittent motion, as shown in Fig. 4, in which the scatter plot of the time $\Delta t_{R}$ for the maximum of $R(t)$ and $\Delta t_{A}$ for the maximum of $A(t)$ is displayed. In almost all analyzed cases, both $\Delta t_{R}$ and $\Delta t_{A}$ were positive, and $\Delta t_{A}$ was slightly greater than $\Delta t_{R}$.
@@
Next, we investigated the dependence of the droplet deformation on the oil red O concentration. We performed experiments at oil red O concentrations in the range of 0.50.50.50.51.$1.1~{}\mathrm{g/L}$. The maximum values of $R(t)$ and $A(t)$ ($R_{\mathrm{max}}$ and $A_{\mathrm{max}}$) increased for oil red O concentrations larger than $0.8~{}\mathrm{g/L}$ as shown in Fig. 5.
@@ III.2 Interfacial tension between oil and water
where $\gamma_{\mathrm{ow}}^{(0)}$ is the interfacial tension at the initial stage, $\Delta\gamma_{\mathrm{ow}}=(\gamma_{\mathrm{ow}}^{(0)}-\gamma_{\mathrm{ow}}(\infty))$ is the difference in interfacial tension between the initial and final stages , and $\tau$ is the characteristic time for the decrease in interfacial tension. This functional form comes from the model describing the rate of relaxation of interfacial tension towards its stable value as proportional to $(\gamma_{\mathrm{ow}}(\infty)-\gamma_{\mathrm{ow}}(t))$. Such behavior can be expected when the dye is absorbed on the droplet surface. We plotted $\gamma_{\mathrm{ow}}^{(0)}$, $\Delta\gamma_{\mathrm{ow}}$ and $\tau$ against oil red O concentration in Fig. 7. $\gamma_{\mathrm{ow}}^{(0)}$ and $\Delta\gamma_{\mathrm{ow}}$ were decreasing and increasing functions of oil red O concentration, respectively. In the case when the decrease in interfacial tension was clearly observed (oil red O concentrations: $0.5$, $0.8$, and $1.0~{}\mathrm{g/L}$), $\tau$ was approximately equal to $800~{}\mathrm{s}$. In addition, we performed the same measurement for the oil droplet that contained both oil red O ($1.0~{}\mathrm{g/L}$) and camphor ($0.06~{}\mathrm{M}$). The decrease in interfacial tension was also observed, and $\gamma_{\mathrm{ow}}^{(0)}$ was lower compared to paraffin droplet with oil red O ($1.0~{}\mathrm{g/L}$) but without camphor.
Figure 6: Time evolution of the interfacial tension $\gamma_{\mathrm{ow}}(t)$ between a paraffin droplet and pure water measured by the pendant drop method. The time $t=0$ corresponds to when a droplet was formed. Marker colors code the oil red O concentrations: $0~{}\mathrm{g/L}$ black, $0.1~{}\mathrm{g/L}$ yellow, $0.5~{}\mathrm{g/L}$ orange, $0.8~{}\mathrm{g/L}$ red, $1.0~{}\mathrm{g/L}$ dark red. For each concentration, the average values of four experiments are plotted. The error bars indicate the standard deviation of the results. $\gamma_{\mathrm{ow}}(t)$ between pure water and paraffin droplet that contained both oil red O ($1.0~{}\mathrm{g/L}$) and camphor ($0.06~{}\mathrm{M}$) is marked by blue squares.
Figure 7: Parameters fitting $\gamma_{\mathrm{ow}}(t)$ according to Eq. (1) for different concentrations of oil red O (cf. Fig. 6). (a) Interfacial tension on the initial stage $\gamma_{\mathrm{ow}}^{(0)}$. (b) Difference in interfacial tension $\Delta\gamma_{\mathrm{ow}}$ between the initial and final stages. (c) Characteristic time $\tau$ for the decrease in interfacial tension. The data for a paraffin droplet that contained both oil red O ($1.0~{}\mathrm{g/L}$) and camphor ($0.06~{}\mathrm{M}$) are marked with blue squares. The average values for four measurements are shown. The error bars indicate their standard deviations.
@@ IV.1 Analysis of free energy related to droplet deformation
To better understand the droplet deformation phenomenon, we evaluated two related dimensionless numbers. First, the capillary number $\mathrm{Ca}$ is the ratio of the viscous force and interfacial tension, $\mathrm{Ca}=\mu v/\gamma_{\mathrm{ow}}$, where $\mu$ is the viscosity of the liquid, $v$ is the characteristic velocity, and $\gamma_{\mathrm{ow}}$ is the characteristic value for oil/water interfacial tension. By substituting $\mu\simeq 2\times 10^{-1}~{}\mathrm{Pa\,s}$ (paraffin), $v\simeq 10^{-2}~{}\mathrm{m/s}$, and $\gamma_{\mathrm{ow}}\simeq 60\times 10^{-3}~{}\mathrm{N/m}$, we obtain $\mathrm{Ca}\simeq 3\times 10^{-2}\ll 1$. This means that during an intermittent motion, the relative effect of viscous drag forces is much smaller than that of surface tension.
Next, the Weber number $\mathrm{We}$ is the relative importance of the fluid inertia compared to its surface tension. $\mathrm{We}=\rho rv^{2}/\gamma_{\mathrm{ow}}$, where $\rho$ is the mass density of the liquid and $r$ is the characteristic length. For $\rho\simeq 10^{3}~{}\mathrm{kg/m^{3}}$, $r\simeq 10^{-2}~{}\mathrm{m}$ and other parameter values listed above, we obtained $\mathrm{We}\simeq 2\times 10^{-5}\ll 1$. This means that the inertial effects are negligible during an intermittent motion. Considering both the dimensionless numbers, we concluded that the surface tension difference around the droplet plays the most important role in droplet deformation.
@@
Here, $S=\gamma_{\mathrm{aw}}-(\gamma_{\mathrm{ow}}+\gamma_{\mathrm{ao}})$ is the spreading coefficient, $\gamma_{\mathrm{aw}}$ and $\gamma_{\mathrm{ao}}$ are the interfacial tensions between air and water and between air and oil, respectively, $g$ is the gravitational acceleration, $\Sigma$ is the area of the droplet, and $\Delta\rho_{\mathrm{ao}}$ and $\Delta\rho_{\mathrm{ow}}$ are the density differences between air and oil and that between oil and water, respectively, as shown in Fig. 8. In Eq. (2), $e_{\mathrm{ao}}$ and $e_{\mathrm{ow}}$ are the differences in height between the water level at infinity and the air/oil interface, and that between the water level at infinity and the oil/water interface, respectively. For the following analyses, we neglected the second term in Eq. (2) as it is much smaller than the first term since the thickness of a droplet is of the order of $100~{}\mu\mathrm{m}$, thus much smaller than the droplet size.
Figure 8: Schematic illustration of an oil droplet floating on the water surface. The triple line matches $x=z=0$. The functions $z=\zeta_{\mathrm{ao}}(x)$ and $z=-\zeta_{\mathrm{ow}}(x)$ indicate the interface between air and oil and that between oil and water, respectively. $\theta_{\mathrm{ao}}(x)$ and $\theta_{\mathrm{ow}}(x)$ determine the slopes of them, i.e., $\tan\theta_{\mathrm{ao}}(x)=\mathrm{d}\zeta_{\mathrm{ao}}/\mathrm{d}x$ and $\tan\theta_{\mathrm{ow}}(x)=\mathrm{d}\zeta_{\mathrm{ow}}/\mathrm{d}x$. $e_{\mathrm{ao}}$ and $e_{\mathrm{ow}}$ are the differences in height between the water level at infinity and the air/oil interface, and that between the water level at infinity and the oil/water interface, respectively.
@@
$\mathcal{T}=$ $\frac{4}{3}\gamma_{\mathrm{ao}}\kappa_{\mathrm{ao}}^{-1}\left[1-\cos^{2}\left(\frac{\theta_{\mathrm{ao}}(x=0)}{2}\right)\right]$ (3)
$+\frac{4}{3}\gamma_{\mathrm{ow}}\kappa_{\mathrm{ow}}^{-1}\left[1-\cos^{2}\left(\frac{\theta_{\mathrm{ow}}(x=0)}{2}\right)\right],$
where $\kappa_{\mathrm{ao}}^{-1}$ and $\kappa_{\mathrm{ow}}^{-1}$ are capillary lengths of air/oil and oil/water interfaces, respectively. $\theta_{\mathrm{ao}}(x=0)$ is the angle between the air/oil interface and the $x$-axis at the triple line, and $\theta_{\mathrm{ow}}(x=0)$ is that between the oil/water interface and the $x$-axis. We calculated the excess energy of the triple line using the relationships linking the angles $\theta_{\mathrm{ao}}(x=0)$, $\theta_{\mathrm{ow}}(x=0)$ with the interfacial tensions derived in Appendix A.3.
@@
To support self-motion with deformation, the energy gain due to the surface tension gradient should exceed the free energy increase due to the droplet deformation depending on the interfacial tension. If the excess energy of the triple line dominates, the system relaxes towards a circular shape as, for example, demonstrated in the case of soap films[45]. In our system, the quasi-elastic laser scattering (QELS) measurements shown in Appendix B suggested that the value of $\delta\gamma_{\mathrm{aw}}$ does not exceed $0.3~{}\mathrm{mN/m}$, though we were not able to measure this value directly. For the following calculations, we used $\delta\gamma_{\mathrm{aw}}=0.25~{}\mathrm{mN/m}$. During the intermittent motion shown in Fig. 3, we observed $\Delta\Sigma=3.3~{}\mathrm{mm}^{2}$, $\Delta l=17.2~{}\mathrm{mm}$ and $\sigma=1288.5~{}\mathrm{mm}^{2}$. For these parameters and values of interfacial tensions measured by the Wilhelmy method: $\gamma_{\mathrm{ao}}=31.6~{}\mathrm{mN/m}$ and $\gamma_{\mathrm{aw}}=73.5~{}\mathrm{mN/m}$, the values of $\Delta E_{\Sigma}=S\Delta\Sigma$, $\Delta E_{l}=\mathcal{T}\Delta l$ and $W$ are plotted as functions of $\gamma_{\mathrm{ow}}$ in Fig. 9(a). For these realistic values of parameters ($\gamma_{\mathrm{ow}}\sim 50~{}\mathrm{mN/m}$), $\Delta E_{\Sigma}+\Delta E_{l}$, and $W$ are of the same order, which implies that the deformation of a droplet can occur. $\Delta E_{\Sigma}(\gamma_{\mathrm{ow}})$ and $\Delta E_{l}(\gamma_{\mathrm{ow}})$ are increasing functions of interfacial tension $\gamma_{\mathrm{ow}}$, while $W$ does not depend on it. The increase in $\Delta E_{l}(\gamma_{\mathrm{ow}})$ is faster than that in $\Delta E_{\Sigma}(\gamma_{\mathrm{ow}})$, which means that the line energy is the most important for droplet deformation. Figure 9(b) shows the deformation energy normalized by the work exerted on the droplet. For the selected parameter values, the spontaneous droplet deformation can occur when $\Delta E/W<1$, which happens if $\gamma_{\mathrm{ow}}\lesssim 47~{}\mathrm{mN/m}.$ Having in mind that $\gamma_{\mathrm{ow}}$ is a decreasing function of oil red O concentration (cf. Fig. 6), we can see why droplet deformation occurs for droplets with a higher concentration of oil red O. The estimated threshold for deformation is a bit lower than expected from the results discussed in the previous chapter. There could be many reasons for this. First, the estimation is based on a simple model. Second, the local inhomogeneities in oil red O concentration, especially those near the droplet boundary, can trigger local shape changes. And finally, there may be possible errors in the measurements of interfacial tension. For example, the threshold should increase for the lower value of $\gamma_{\mathrm{ao}}$.
@@
Figure 9: Major forms of energy related to droplet deformation as functions of the interfacial tension between oil and water. (a) Plots of $\Delta E_{\Sigma}$ (red), $\Delta E_{l}$ (orange) and $W$ (blue), respectively. (b) Solid line shows deformation energy normalized by the work exerted on the droplet, $\Delta E/W$ against $\gamma_{\mathrm{ow}}$.